Calculate the theoretical yield of iron(III) oxide (Fe2O3) in grams.


1
Expert's answer
2021-12-10T03:26:46-0500

The reaction is 2 Fe + O2   -------  2 FeO

4.8 g O2 = 4.8/32 = 0.15 mol.

0.15 mol O2 produces 2 x 0.15 = 0.3 mol FeO. Multiplying this by the molar mass of FeO (72) produces 21.6 grams.

Theoretical yield = 21.6 grams

% yield = 18.4/21.6 = 85%